Could you help with this problem:
1/3S + 7/9 = 4/3s
So far what I have done is:
Multiplied both sides by 9 to get 9(1)/3s + 9(7)/9 + 9(4)3s
Then I get 3 * s + 7 = 3 * 4 * s
Which is 3 * s + 7 = 12s
Then I have subtracted 3s from both sides but I'm not sure about these steps.
I get stuck here. Can you help?

Answer by stanbon(75887)

Then I get 3 * s + 7 = 3 * 4 * s
Which is 3 * s + 7 = 12s
Then subtract 3s from both sides to get:
7=9s
Divide both sides by 9 to get
x=7/9
Cheers,
Stan H.